# Mixture of Gaussian-distributed Prototypes with Generative Modelling for Interpretable and Trustworthy Image Recognition

### v2

@@ Abstract
Prototypical-part methods, e.g., ProtoPNet, enhance interpretability in image recognition by linking predictions to training prototypes, thereby offering intuitive insights into their decision-making. Existing methods, which rely on a point-based learning of prototypes, typically face two critical issues: 1) the learned prototypes have limited representation power and are not suitable to detect Out-of-Distribution (OoD) inputs, reducing their decision trustworthiness; and 2) the necessary projection of the learned prototypes back into the space of training images causes a drastic degradation in the predictive performance. Furthermore, current prototype learning adopts an aggressive approach that considers only the most active object parts during training, while overlooking sub-salient object regions which still hold crucial classification information. In this paper, we present a new generative paradigm to learn prototype distributions, termed as Mixture of Gaussian-distributed Prototypes (MGProto). The distribution of prototypes from MGProto enables both interpretable image classification and trustworthy recognition of OoD inputs. The optimisation of MGProto naturally projects the learned prototype distributions back into the training image space, thereby addressing the performance degradation caused by prototype projection. Additionally, we develop a novel and effective prototype mining strategy that considers not only the most active but also sub-salient object parts. To promote model compactness, we further propose to prune MGProto by removing prototypes with low importance priors. Experiments on CUB-200-2011, Stanford Cars, Stanford Dogs, and Oxford-IIIT Pets datasets show that MGProto achieves state-of-the-art image recognition and OoD detection performances, while providing encouraging interpretability results.
@@
Interpretability, prototypical-part networks, Gaussian mixture models, generative modelling, horse racing, prototype mining.
I Introduction
Deep learning models [1, 2] show remarkable performance in computer vision tasks, but their complex network architectures and massive parameter spaces make their decision processes opaque [3, 4]. As a result, deep learning models may not be trustworthy, particularly in safety-critical domains where the consequences of model errors can be severe [4, 5, 6, 7]. To alleviate this issue, explainable artificial intelligence (XAI) has gained increasing focus, striving to develop interpretable strategies to explain the internal workings of deep learning models in a manner comprehensible to humans [8, 9, 10, 11, 12]. Among them, the prototypical-part network (ProtoPNet) [10] is an appealing grey-box approach that classifies images based on similarities of the test sample to the prototypes corresponding to image patches previously-seen in the training set. Such strategy relies on train-test sample associations for decision making, which resembles how humans reason according to cognitive psychological studies [13, 14] revealing that humans use past cases as models when learning to solve problems. The success of ProtoPNet has motivated the development of its many variants, e.g., TesNet [15], ProtoTree [16], and PIP-Net [17].
Current prototypical part networks [10, 17, 18] depend on a discriminative classifier [19] optimised with point-based learning techniques that train prototypes with specific values. Such a classifier produces logits (computed from a weighted sum of similarity scores), which are passed on to a softmax function to directly obtain the class probability given an input image, i.e., $p(c|\mathbf{x})$, as shown in Fig. 1(a). Though straightforward and easy to implement, the resulting prototypes from point-based learning have limited representation power and tends to suffer from drastic performance degradation (see Fig. 2) upon the prototype replacement [^1] In the literature, the prototype replacement is also referred to as prototype “pushing” or “projection”. In this paper, we will use the terms replacement and projection interchangeably. that is an indispensable step for grounding prototypes in the human-understandable space of training images [20]. Unfortunately, current methods either omit this replacement step [21, 22, 23, 24] for better task accuracy, but at the cost of compromising interpretability given that their prototypes are no longer represented by actual training image patches, or include the replacement step without addressing the performance degradation issue [10, 15, 25, 18]. Additionally, despite producing encouraging interpretable results, these discriminative classifiers miss an important explainability feature [26, 27], the recognition of Out-of-Distribution (OoD) inputs. Ideally, an XAI system should be able to explain its predictions regarding In-Distribution (ID) samples (interpretable) and meanwhile identify anomalous OoD samples to abstain from classifying them to ensure the decision trustworthiness. Furthermore, the prototype learning in current methods is an aggressive approach that considers only the most salient object parts, disregarding important information available from less-salient object regions which could be helpful in the encoding of more difficult-to-learn visual features that have the potential to achieve improved classification.
@@ I Introduction
II Related Work
II-A Interpretability with Prototypes
Efforts to understand deep learning networks often face challenges if relying solely on post-hoc techniques like saliency maps [30, 31, 32], explanation surrogates [33, 34, 35], and counterfactual examples [36, 37, 38]. These strategies fall short in explaining the network’s reasoning process, yielding results that may be unreliable and risky [3].
On the other hand, prototype-based networks provide an appealing way to access the model’s inner workings. These networks dissect a test image sample by finding prototypes and combining evidence from these prototypes to make a decision. ProtoPNet [10] is the original work that adopts a number of class-specific prototypes for interpretable image recognition. Built upon it, TesNet [15] introduces the orthonormal transparent basis concepts, and Deformable ProtoPNet [25] makes prototypes spatially-deformable to handle object variations. Other efforts extend prototypes to visual transformers [22, 24], K-nearest neighbour [39], knowledge distillation [40], and ensemble interpretations [18]. To gauge the prototype-based interpretability, two objective metrics are proposed, i.e., consistency and stability scores [21]. Also, some studies relax the class-specific constraint to allow a reduced number of prototypes [41, 42, 16, 24]. Recent works focus on refining semantically confounded or ambiguous prototypes [43, 44] in ProtoPNet. Promisingly, the prototype-based interpretable networks are being applied in critical tasks beyond the ones in computer vision [45, 46, 47, 48].
The existing models above are trained with point-based learning techniques, yielding prototypes with limited representation power. As described in Section I, the training for these prototypes involves a prototype projection into the training image space that causes drastic performance drop. Furthermore, it is challenging to recognise OoD inputs using these prototypes trained with point-based learning methods, thereby reducing the decision trustworthiness of the model. Even though PIP-Net [17], trained with point-based learning techniques, attempts to detect OoD samples, its detection performance is still far from satisfactory.
II-B Gaussian Mixture Model
@@
II-C Tian Ji’s Horse Racing Legend
@@
The mathematical principle of this victorious strategy achieved more than 2000 years ago is still studied in modern game theory [56, 57]. In this work, we take inspiration from this legend to build an infallible winning strategy for the King: if the regular and the slow horses of the King are trained to be faster than the fast horse of Tian, then the King will always win the race of horse competing at each round. To the best of our knowledge, we are the first to employ the horse-racing strategy in XAI research.
III Our MGProto Method
This section starts with a brief description of the current ProtoPNet-based networks with point-based learning of prototypes in Section III-A, serving as preliminaries for our proposed MGProto method illustrated in Fig. 4. Then, we detail our MGProto that represents prototypes with Gaussian mixture models (Section III-B), which are learned via a modified EM algorithm to ensure prototype diversity (Section III-C). Additionally, we present a novel and generic approach to mine prototypes from sub-salient object regions to improve classification, as elaborated in Section III-D.
III-A Preliminaries
@@
III-B Mixture of Gaussian-distributed Prototypes
According to Section III-A, current discriminative networks rely on point-based learned prototypes $\mathcal{P}$ that are used to compute similarity maps with $\mathbf{F}$. Instead, we leverage a Gaussian distribution to model our distribution-based prototypes and obtain likelihood maps whose value at the spatial position $(i,j)$ denotes the confidence that the image patch fits the $m$-th prototype distribution of class $c$, as in:
@@ III-B Mixture of Gaussian-distributed Prototypes
where $\boldsymbol{\pi}_{m}^{c}$ is the mixture weights in GMMs to quantify prototype importance, which are refered to as importance priors. Also, $\boldsymbol{\pi}_{m}^{c}$ serves as the weights in the linear aggregating layer. From Eq. (3), the mixture model allows the prototype distributions of a class to collaboratively describe the underlying data density of that class. Meanwhile, these multiple prototype distributions in a mixture are enforced to compete against each other due to the inherent constraint in GMM: $\sum_{m=1}^{M}\boldsymbol{\pi}_{m}^{c}=1$ for each class $c$. Hence, we take advantage of this constraint to prune prototypes with low importance priors for promoting model compactness. It is worth noting that our MGProto utilises a concise aggregating layer (see Fig. 1(b)), represented by $\boldsymbol{\pi}^{c}\in\mathbb{R}^{M\times 1}$, which differs from current methods [10, 25, 18] using a dense layer denoted by $\boldsymbol{\pi}^{c}\in\mathbb{R}^{MC\times 1}$, as mentioned in Section III-A. The use of our concise layer not only introduces prototype-class connection sparsity to reduce the explanation size [17], but also prevents the prediction of a class from being disturbed by prototypes of other classes [21].
@@
by assuming $p(c)=\frac{1}{C}$, as commonly adopted in the literature [58, 59, 60]. Relying on this assumption, the core of MGProto is to accurately estimate the class-conditional data density $p(\mathbf{x}|c)$ modelled with the mixture of Gaussian-distributed prototypes through Eq. (3).
@@
The well-known expectation-maximisation (EM) algorithm [55] can be employed to estimate the prototype means $\mathbf{p}^{c}$ and importance priors $\boldsymbol{\pi}^{c}$ for each class $c$. Nevertheless, the standard EM algorithm does not guarantee diverse characteristics of the Gaussian-distributed prototypes, a crucial aspect for enhancing the interpretability of prototypical-part methods [15, 25, 18]. Therefore, in the next Section III-C, we introduce a modified EM strategy to encourage diversity of the prototypes during their learning.
III-C Learning of Gaussian-distributed Prototypes
Section III-B describes the prototype means $\mathbf{p}^{c}$ and importance priors $\boldsymbol{\pi}^{c}$ as parameters of the class-wise GMMs, defined in Eq. (3). The accurate optimisation of GMM parameters typically depends on the whole training set [49]. However, modern deep-learning networks are often trained in a mini-batch fashion, where only a limited number of training data is provided. That can harm GMM’s predicative performance. To leverage a large set of feature representations for the GMM optimisation, we adopt the memory bank [62] mechanism that has proven effective in maintaining contextual information [63].
@@ III-C Learning of Gaussian-distributed Prototypes
III-D Mining of Prototypes from Sub-salient Object Parts
@@
We elaborate below how to formulate the new prototype mining strategy. Assume that we have a training image $\mathbf{x}$ labelled with class $c$. We consider $T$ sequential patch levels from the most to the least active, such as $\{A^{1},A^{2},...,A^{T}\}$ of the prototype $\mathbf{p}_{1}$ and $\{B^{1},B^{2},...,B^{T}\}$ of the prototype $\mathbf{p}_{2}$, as in Fig. 5(a). According to Eq. (3), the predicted logit [^2] In our generative MGProto, the class-conditional data probability $p(\mathbf{x}|c)$ serves as the model’s prediction logits, but we still use the symbol $\mathbf{logit}$ here, given the generality of the proposed mining strategy to the existing discriminative-based prototypical-part networks. for each patch level $t\in\{1,...,T\}$ is computed as the weighted sum of likelihood scores $p(\mathbf{x}|c)$: $\mathbf{logit}^{c}_{t}=\boldsymbol{\pi}_{1}^{c}\times\mathbf{H}_{1}^{(A^{t}),c}+\boldsymbol{\pi}_{2}^{c}\times\mathbf{H}_{1}^{(B^{t}),c}$. Obviously, we have $\mathbf{logit}^{c}_{T}<\mathbf{logit}^{c}_{T-1}<...<\mathbf{logit}^{c}_{1}$ because $\mathbf{H}_{1}^{(A^{T}),c}<\mathbf{H}_{1}^{(A^{T-1}),c}<...<\mathbf{H}_{1}^{(A^{1}),c}$ and $\mathbf{H}_{1}^{(B^{T}),c}<\mathbf{H}_{1}^{(B^{T-1}),c}<...<\mathbf{H}_{1}^{(B^{1}),c}$. We showcase our prototype mining strategy in Fig. 5(c), which is analogous to the legend. To propose the infallible winning strategy for the King, we regard: 1) the most active logit of the wrong classes $\lnot c$, denoted by $\mathbf{logit}^{\lnot c}_{1}\in\mathbb{R}^{C-1}$, as Tian Ji’s fast horse; and 2) the $T-1$ less active logits of the actual class $c$, denoted by $\mathbf{logit}^{c}_{t\in\{2,...,T\}}$, as the King’s horses, except for the most active one denoted by $\mathbf{logit}^{c}_{1}$. Then, we devise $T-1$ mining competitions: $\mathbf{logit}^{\lnot c}_{1}$ vs. $\mathbf{logit}^{c}_{t\in\{2,...,T\}}$, which are supervised by the following mining loss with the goal of increasing $\mathbf{logit}^{c}_{t\in\{2,...,T\}}$:
$\mathcal{L}_{mn}(\mathbf{x},\mathbf{y})=\frac{1}{T-1}\sum_{t=2}^{T}\mathcal{L}_{ce}([\mathbf{logit}^{\lnot c}_{1},\mathbf{logit}^{c}_{t}],\mathbf{y}),$ (10)
@@
where $\mathcal{L}_{ce}(\cdot)$ is the cross-entropy (CE) loss and $\mathbf{y}$ is the ground-truth label of $\mathbf{x}$. Here we employ $[\mathbf{logit}^{\lnot c}_{1},\mathbf{logit}^{c}_{t}]\in\mathbb{R}^{C}$ to represent the combined mining logits corresponding to the solid lines in Fig. 5(c), with $[\mathbf{logit}^{\lnot c}_{1},\mathbf{logit}^{c}_{1}]\in\mathbb{R}^{C}$ denoting the standard classification logits, shown as the dash line, which is supervised by:
$\mathcal{L}_{ce}(\mathbf{x},\mathbf{y})=\mathcal{L}_{ce}([\mathbf{logit}^{\lnot c}_{1},\mathbf{logit}^{c}_{1}],\mathbf{y}),$ (11)
@@
Different from the legend in Fig. 5(b), our prototype mining leverages all $\mathbf{logit}^{c}_{t\in\{2,...,T\}}$ to compete with the unique and most active $\mathbf{logit}^{\lnot c}_{1}$ to ensure an infallible winning strategy for the King, which also yields a stronger mining effect. Our proposed prototype mining strategy can be readily applied to a wide range of prototypical-part networks, for improved classification.
III-E Training Objective and Prototype Replacement

### v3

@@ Abstract
Prototypical-part methods, e.g., ProtoPNet, enhance interpretability in image recognition by linking predictions to training prototypes, thereby offering intuitive insights into their decision-making. Existing methods, which rely on a point-based learning of prototypes, typically face two critical issues: 1) the learned prototypes have limited representation power and are not suitable to detect Out-of-Distribution (OoD) inputs, reducing their decision trustworthiness; and 2) the necessary projection of the learned prototypes back into the space of training images causes a drastic degradation in the predictive performance. Furthermore, current prototype learning adopts an aggressive approach that considers only the most active object parts during training, while overlooking sub-salient object regions which still hold crucial classification information. In this paper, we present a new generative paradigm to learn prototype distributions, termed as Mixture of Gaussian-distributed Prototypes (MGProto). The distribution of prototypes from MGProto enables both interpretable image classification and trustworthy recognition of OoD inputs. The optimisation of MGProto naturally projects the learned prototype distributions back into the training image space, thereby addressing the performance degradation caused by prototype projection. Additionally, we develop a novel and effective prototype mining strategy that considers not only the most active but also sub-salient object parts. To promote model compactness, we further propose to prune MGProto by removing prototypes with low importance priors. Experiments on CUB-200-2011, Stanford Cars, Stanford Dogs, and Oxford-IIIT Pets datasets show that MGProto achieves state-of-the-art image recognition and OoD detection performances, while providing encouraging interpretability results. Code is available at https://github.com/cwangrun/MGProto.
@@
Interpretability, prototypical-part networks, Gaussian mixture, horse racing, generative modelling, prototype mining.
1 Introduction
Deep learning models [1, 2] show remarkable performance in computer vision tasks, but their complex network architectures, high non-linearity, and massive parameter spaces make their decision processes opaque [3, 4]. As a result, deep learning models may not be trustworthy, particularly in safety-critical domains where the consequences of model errors can be severe [4, 5, 6, 7]. To alleviate this issue, explainable artificial intelligence (XAI) has gained increasing focus, striving to develop interpretable strategies to explain the internal workings of deep learning models in a manner comprehensible to humans [8, 9, 10, 11, 12]. Among them, the prototypical-part network (ProtoPNet) [10] is an appealing grey-box approach that classifies images based on similarities of the test sample to the prototypes corresponding to image patches previously-seen in the training set. Such strategy relies on train-test sample associations for decision making, which resembles how humans reason according to cognitive psychological studies [13, 14] revealing that humans use past cases as models when learning to solve problems. The success of ProtoPNet has motivated the development of its many variants, e.g., TesNet [15], ProtoTree [16], and PIP-Net [17].
@@ 1 Introduction
Current prototypical part networks [10, 17, 18] depend on a discriminative classifier [19] optimised with point-based learning techniques that train prototypes with specific values. Such a classifier produces logits (computed from a weighted sum of similarity scores), which are passed on to a softmax function to directly obtain the class probability given an input image, i.e., $p(c|\mathbf{x})$, as shown in Fig. 1(a). Though straightforward and easy to implement, the resulting prototypes from point-based learning have limited representation power and tends to suffer from drastic performance degradation (see Fig. 2) upon the prototype replacement [^1] In the literature, the prototype replacement is also referred to as prototype “pushing” or “projection”. In this paper, we will use the terms replacement and projection interchangeably. that is an indispensable step for grounding prototypes in the human-understandable space of training images [20]. Unfortunately, current methods either omit this replacement step [21, 22, 23, 24] for better task accuracy, but at the cost of compromising interpretability given that their prototypes are no longer represented by actual training image patches, or include the replacement step without addressing the performance degradation issue [10, 15, 25, 18]. Additionally, despite producing encouraging interpretable results, these discriminative classifiers miss an important explainability feature [26, 27], the recognition of Out-of-Distribution (OoD) inputs. Ideally, an XAI system should be able to explain its predictions regarding In-Distribution (ID) samples (interpretable) and meanwhile identify anomalous OoD samples to abstain from classifying them to ensure the decision trustworthiness. Furthermore, the prototype learning in current methods is an aggressive222“Aggressive” refers to the fact that the prototype’s training process relies exclusively on the most salient object regions. approach that considers only the most salient object parts, disregarding important information available from less-salient object regions which could be helpful in the encoding of more difficult-to-learn visual features that have the potential to achieve improved classification.
@@
2 Related Work
2.1 Interpretability with Prototypes
Efforts to understand deep learning networks often face challenges if relying solely on post-hoc techniques like saliency maps [30, 31, 32, 33], explanation surrogates [34, 35, 36], and counterfactual examples [37, 38, 39]. These strategies fall short in explaining the network’s reasoning process, yielding results that may be unreliable and risky [3].
On the other hand, prototype-based networks provide an appealing way to access the model’s inner workings. These networks dissect a test image sample by finding prototypes and combining evidence from these prototypes to make a decision. ProtoPNet [10] is the original work that adopts a number of class-specific prototypes for interpretable image recognition. Built upon it, TesNet [15] introduces the orthonormal transparent basis concepts, and Deformable ProtoPNet [25] makes prototypes spatially-deformable to handle object variations. Other efforts extend prototypes to visual transformers [22, 24], K-nearest neighbour [40], knowledge distillation [41], and ensemble interpretations [18]. To gauge the prototype-based interpretability, two objective metrics are proposed, i.e., consistency and stability scores [21]. Also, some studies relax the class-specific constraint to allow a reduced number of prototypes [42, 43, 16, 24]. Recent works focus on refining semantically confounded or ambiguous prototypes [44, 45] in ProtoPNet. Promisingly, the prototype-based interpretable networks are being applied in critical tasks beyond the ones in computer vision [46, 47, 48, 49, 50].
The existing models above are trained with point-based learning techniques, yielding prototypes with limited representation power. As described in Section 1, the training for these prototypes involves a prototype projection into the training image space that causes drastic performance drop. Furthermore, it is challenging to recognise OoD inputs using these prototypes trained with point-based learning methods, thereby reducing the decision trustworthiness of the model. Even though PIP-Net [17], trained with point-based learning techniques, attempts to detect OoD samples, its detection performance is still far from satisfactory.
2.2 Gaussian Mixture Model
@@
2.3 Tian Ji’s Horse Racing Legend
@@
The mathematical principle of this victorious strategy achieved more than 2000 years ago is still studied in modern game theory [58, 59]. In this work, we take inspiration from this legend to build an infallible winning strategy for the King: if the regular and the slow horses of the King are trained to be faster than the Tian Ji’s fast horse, then the King will always win the race by winning at each round. This infallible winning strategy of the King resembles our mining of prototypes from sub-salient object parts, elaborated in Section 3.4. To the best of our knowledge, we are the first to employ the horse-racing strategy in XAI research.
3 Our MGProto Method
This section starts with a brief description of the current ProtoPNet-based networks with point-based learning of prototypes in Section 3.1, serving as preliminaries for our proposed MGProto method illustrated in Fig. 4. Then, we detail our MGProto that represents prototypes with Gaussian mixture models (Section 3.2), which are learned via a modified EM algorithm to ensure prototype diversity (Section 3.3). Additionally, we present a novel and generic approach to mine prototypes from sub-salient object regions to improve classification, as elaborated in Section 3.4.
3.1 Preliminaries
@@
3.2 Mixture of Gaussian-distributed Prototypes
According to Section 3.1, current discriminative networks rely on point-based learned prototypes $\mathcal{P}$ that are used to compute similarity maps with $\mathbf{F}$. Instead, we leverage a Gaussian distribution to model our distribution-based prototypes and obtain likelihood maps whose value at the spatial position $(i,j)$ denotes the confidence that the image patch fits the $m$-th prototype distribution of class $c$, as in:
@@ 3.2 Mixture of Gaussian-distributed Prototypes
where $\boldsymbol{\pi}_{m}^{c}$ is the mixture weights in GMMs to quantify prototype importance, which are refered to as importance priors. Also, $\boldsymbol{\pi}_{m}^{c}$ serves as the weights in the linear aggregating layer. From Eq. (3), the mixture model allows the prototype distributions of a class to collaboratively describe the underlying data density of that class. Meanwhile, these multiple prototype distributions in a mixture are enforced to compete against each other due to the inherent constraint in GMM: $\sum_{m=1}^{M}\boldsymbol{\pi}_{m}^{c}=1$ for each class $c$. Hence, we take advantage of this constraint to prune prototypes with low importance priors for promoting model compactness. It is worth noting that our MGProto utilises a concise aggregating layer (see Fig. 1(b)), represented by $\boldsymbol{\pi}^{c}\in\mathbb{R}^{M\times 1}$, which differs from current methods [10, 25, 18] using a dense layer denoted by $\boldsymbol{\pi}^{c}\in\mathbb{R}^{MC\times 1}$, as mentioned in Section 3.1. The use of our concise layer not only introduces prototype-class connection sparsity to reduce the explanation size [17], but also prevents the prediction of a class from being disturbed by prototypes of other classes [21].
@@
by assuming $p(c)=\frac{1}{C}$ [^3] The datasets used in our work have fairly balanced training samples across classes, making the prior $p(c)$ identical for all classes. For the imbalanced scenario, the prior can be estimated from the observed training data distribution. Alternatively, resampling strategies [60] can be used during training, properly ensuring an identical prior., as commonly adopted in the literature [61, 62, 63]. Relying on this assumption, the core of MGProto is to accurately estimate the class-conditional data density $p(\mathbf{x}|c)$ modelled with the mixture of Gaussian-distributed prototypes through Eq. (3).
@@
The well-known expectation-maximisation (EM) algorithm [57] can be employed to estimate the prototype means $\mathbf{p}^{c}$ and importance priors $\boldsymbol{\pi}^{c}$ for each class $c$. Nevertheless, the standard EM algorithm does not guarantee diverse characteristics of the Gaussian-distributed prototypes, a crucial aspect for enhancing the interpretability of prototypical-part methods [15, 25, 18]. Therefore, in the next Section 3.3, we introduce a modified EM strategy to encourage diversity of the prototypes during their learning.
3.3 Learning of Gaussian-distributed Prototypes
Section 3.2 describes the prototype means $\mathbf{p}^{c}$ and importance priors $\boldsymbol{\pi}^{c}$ as parameters of the class-wise GMMs, defined in Eq. (3). The accurate optimisation of GMM parameters typically depends on the whole training set [51]. However, modern deep-learning networks are often trained in a mini-batch fashion, where only a limited number of training data is provided. That can harm GMM’s predicative performance. To leverage a large set of feature representations for the GMM optimisation, we adopt the memory bank [65] mechanism that has proven effective in maintaining contextual information [66].
@@ 3.3 Learning of Gaussian-distributed Prototypes
3.4 Mining of Prototypes from Sub-salient Object Parts
@@
We elaborate below how to formulate the new prototype mining strategy. Assume that we have a training image $\mathbf{x}$ labelled with class $c$. We consider $T$ sequential patch levels from the most to the least active, such as $\{A^{1},A^{2},...,A^{T}\}$ of the prototype $\mathbf{p}_{1}$ and $\{B^{1},B^{2},...,B^{T}\}$ of the prototype $\mathbf{p}_{2}$, as in Fig. 5(a). According to Eq. (3), the predicted logit [^4] In our generative MGProto, the class-conditional data probability $p(\mathbf{x}|c)$ serves as the model’s prediction logits, but we still use the symbol $\mathbf{logit}$ here, given the generality of the proposed mining strategy to the existing discriminative-based prototypical-part networks. for each patch level $t\in\{1,...,T\}$ is computed as the weighted sum of likelihood scores $p(\mathbf{x}|c)$: $\mathbf{logit}^{c}_{t}=\boldsymbol{\pi}_{1}^{c}\times\mathbf{H}_{1}^{(A^{t}),c}+\boldsymbol{\pi}_{2}^{c}\times\mathbf{H}_{1}^{(B^{t}),c}$. Obviously, we have $\mathbf{logit}^{c}_{T}<\mathbf{logit}^{c}_{T-1}<...<\mathbf{logit}^{c}_{1}$ because $\mathbf{H}_{1}^{(A^{T}),c}<\mathbf{H}_{1}^{(A^{T-1}),c}<...<\mathbf{H}_{1}^{(A^{1}),c}$ and $\mathbf{H}_{1}^{(B^{T}),c}<\mathbf{H}_{1}^{(B^{T-1}),c}<...<\mathbf{H}_{1}^{(B^{1}),c}$. We showcase our prototype mining strategy in Fig. 5(c), which is analogous to the legend. To propose the infallible winning strategy for the King, we regard: 1) the most active logit of the wrong classes $\lnot c$, denoted by $\mathbf{logits}^{\lnot c}_{1}\in\mathbb{R}^{C-1}$, as Tian Ji’s fast horse; and 2) the $T-1$ less active logits of the actual class $c$, denoted by $\mathbf{logit}^{c}_{t\in\{2,...,T\}}$, as the King’s horses, except for the most active one denoted by $\mathbf{logit}^{c}_{1}$. Then, we devise $T-1$ mining competitions: $\mathbf{logits}^{\lnot c}_{1}$ vs. $\mathbf{logit}^{c}_{t\in\{2,...,T\}}$, which are supervised by the following mining loss with the goal of increasing $\mathbf{logit}^{c}_{t\in\{2,...,T\}}$:
$\mathcal{L}_{mn}(\mathbf{x},\mathbf{y})=\frac{1}{T-1}\sum_{t=2}^{T}\mathcal{L}_{ce}([\mathbf{logits}^{\lnot c}_{1},\mathbf{logit}^{c}_{t}],\mathbf{y}),$ (10)
@@
where $\mathcal{L}_{ce}(\cdot)$ is the cross-entropy (CE) loss and $\mathbf{y}$ is the ground-truth label of $\mathbf{x}$. Here we employ $[\mathbf{logits}^{\lnot c}_{1},\mathbf{logit}^{c}_{t}]\in\mathbb{R}^{C}$ to represent the combined mining logits corresponding to the solid lines in Fig. 5(c), with $[\mathbf{logits}^{\lnot c}_{1},\mathbf{logit}^{c}_{1}]\in\mathbb{R}^{C}$ denoting the standard classification logits, shown as the dash line, which is supervised by:
$\mathcal{L}_{ce}(\mathbf{x},\mathbf{y})=\mathcal{L}_{ce}([\mathbf{logits}^{\lnot c}_{1},\mathbf{logit}^{c}_{1}],\mathbf{y}),$ (11)
@@
Different from the legend in Fig. 5(b), our prototype mining leverages all $\mathbf{logit}^{c}_{t\in\{2,...,T\}}$ with $T-1$ patch levels as in Eq. (10) to compete with the unique and most active $\mathbf{logits}^{\lnot c}_{1}$ to ensure an infallible winning strategy for the King, which also yields a stronger mining effect. To sum up, the major connections between our prototype mining and the classic horse-racing legend are: Tian Ji achieves victory through clever tactics in the legend. In contrast, our prototype mining approach devises a guaranteed winning strategy for the King, by training the King’s regular and slower horses to ensure they surpass Tian Ji’s fastest horse. Our proposed prototype mining strategy can be readily applied to a wide range of prototypical-part networks, for improved classification.
3.5 Training Objective and Prototype Replacement